OpenStudy (anonymous):

Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (anonymous):

yikes

OpenStudy (anonymous):

Lol, oh no.

OpenStudy (anonymous):

is it \(\tan^3(x)\) (i hope) or \(\tan(3x)\) (i hope not)

OpenStudy (anonymous):

one minute.

OpenStudy (anonymous):

1st

OpenStudy (anonymous):

whew

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

ok lets work on the right only, it is mostly algebra and to make it easier, lets replace cosine by \(a\) and sine by \(b\)

OpenStudy (anonymous):

you have with those replacements \[\frac{a}{b}+\frac{2b}{a}+\frac{b^2}{a^3}\] to add those fractions we need the common denominator which is \(a^3b\)

OpenStudy (anonymous):

oops typo there \[\frac{a}{b}+\frac{2b}{a}+\frac{b^3}{a^3}\]

OpenStudy (anonymous):

add just like always \[\frac{a^4+2a^2b^2+b^4}{a^3b}\]

OpenStudy (anonymous):

you with me so far?

OpenStudy (anonymous):

yea, so

OpenStudy (anonymous):

don't you subtract your denominator & numerator.

OpenStudy (anonymous):

you said add them?

OpenStudy (anonymous):

no it is addition only

OpenStudy (anonymous):

a^7 + 2a^5b^3 + b^5

OpenStudy (anonymous):

lets go slow

OpenStudy (anonymous):

something like that?

OpenStudy (anonymous):

you good to here \[\frac{a}{b}+\frac{2b}{a}+\frac{b^3}{a^3}\]?

OpenStudy (anonymous):

this is where you say "yes, i get that part"

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok now the denomiators are \(a,b,a^3\) the least common multiply of those three is \(a^3b\)

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{2b}{a}+\frac{b^3}{a^3}\] \[=\frac{a}{b}\times \frac{a^3}{a^3}+\frac{2b}{a}\times \frac{a^2b}{a^2b}+\frac{b^3}{a^3}\times \frac{b}{b}\]\[=\frac{a^4+2a^2b^2+b^4}{a^3b}\]

OpenStudy (anonymous):

one more step the numerator is a perfect square it is \[(a^2+b^2)^2\]

OpenStudy (anonymous):

now back to trig since \(a=\cos(x), b =\sin(x)\)you have \[\frac{(\cos^2(x)+\sin^2(x))^2}{\cos^3(x)\sin(x)}\]

OpenStudy (anonymous):

and we recall that \[\sin^2(x)+\cos^2(x)=1\] so the numerator is only a one \[\frac{(\cos^2(x)+\sin^2(x))^2}{\cos^3(x)\sin(x)}=\frac{1}{\cos^3(x)\sin(x)}\]

OpenStudy (anonymous):

From the left hand side: sec^2 = 1+tan^2, hence, sec^4 = (1+tan^2)^2therefore, sec^4 = 1+ 2 tan^2 + tan^4 replace allcot ( 1+ 2 tan^2 + tan^4) distribute cot inside the bracket, cancel more by using cot = 1tanto get the right hand side

OpenStudy (anonymous):

i was working from the right to the left you can work the other way if you like but i find it more confusing

OpenStudy (anonymous):

besides it pretty much says to work from right to left if i read it correctly

OpenStudy (anonymous):

Oh right, ok. continue :)

OpenStudy (anonymous):

\[\frac{1}{\cos^3(x)\sin(x)}=\frac{\cos(x)}{\cos^4(x)\sin(x)}=\frac{1}{\cos^4(x)}\times \frac{\cos(x)}{\sin(x)}=\sec^4(x)\cot(x)\]

OpenStudy (anonymous):

whew that was exhauting now i have to go have a drink

OpenStudy (anonymous):

That's it?

OpenStudy (anonymous):

that is it yes

OpenStudy (anonymous):

My teacher's insane...

OpenStudy (anonymous):

teaching math will do that for you

OpenStudy (anonymous):

go back over the steps, make sure you understand them all steps are there i have to run, good luck

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

I will...

OpenStudy (anonymous):

btw convince yourself that it is 95% algebra, very little trig yw

OpenStudy (anonymous):

Ok :)